# On Glashier’s Partition Theorem

### v1

@@ Abstract.
Glashier’s theorem states that the number of partitions of $n$ into parts which repeat at most m1m-1 times is equal to the number of partitions of $n$ into parts which are not divisible by $m$. The $m=2$ case is Euler’s famous partition theorem. Recently, Andrews, Kumar, and Yee gave two new partition functions $C(n)$ and $D(n)$ related to Euler’s theorem. Lin and Zhang extended their result to Glashier’s theorem by generalizing $C(n)$. We generalize $D(n)$, prove an analogous partition identity for the $m=3$ case, and provide a new series equal to Glashier’s product both in the finite and infinite cases.
@@
Partitions, Partition Identities, Euler’s Theorem, Glashier’s Theorem, $q$-series.
@@ 1. Introduction
([3, Theorem $1.1$]) For any positive integer $n$, we have
@@
Let $m\geq 2$ be an integer. Define $A_{m}(n)$ to be the number of partitions of n into parts repeating less than $m$ times and $B_{m}(n)$ to be the number of partitions of $n$ into parts not divisible by $m$. In 1883, Glashier [4] obtained the following partition theorem which is a natural extension of Euler’s theorem (1.1).
@@
([5, Theorem $1.5$]) For any positive integer $n$, we have
@@
(1.7) $\sum\limits_{n=0}^{\infty}E_{m}(n)q^{n}=\sum\limits_{n=0}^{\infty}q^{mn}(q^{n+1};q)_{\infty}\sum\limits_{j=1}^{m-1}(\zeta^{j}_{m}q^{n+1};q)_{\infty}.$
@@
(1.8) $\varepsilon(q):=\sum\limits_{n=0}^{\infty}E_{3}(n)q^{n}=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}+1},$
where $\chi(n)$ is defined as
$\chi(n):=\omega^{n}+\omega^{-n}=\Bigg\{\begin{array}[]{lr}2\qquad\,\,\,\text{if }3\mid n,\\ -1\quad\,\,\,\,\text{if }3\nmid n.\end{array}$
@@
In view of the generating function of $B_{2}(n)$ in [3, $(2.1)$], we also obtain the following finite $q$-series identity.
@@
The rest of the paper is organized as follows. In Sections 2 and 3, we provide $q$-series (generating function) proofs of Theorem 1.4 and 1.6 respectively.
@@
(3.1) $\varepsilon(q)=(q;q)_{\infty}\sum\limits_{j=0}^{\infty}\dfrac{q^{3j}}{(q;q)_{j}}\left((\omega q^{j+1};q)_{\infty}+(\omega^{-1}q^{j+1};q)_{\infty}\right).$
Now, define $\chi(n)$ as follows
$\chi(n):=\omega^{n}+\omega^{-n}=\Bigg\{\begin{array}[]{lr}2\qquad\,\,\,\text{if }3\mid n,\\ -1\quad\,\,\,\,\text{if }3\nmid n.\end{array}$
@@
(3.2) $\varepsilon(q)=(q;q)_{\infty}\sum\limits_{j=0}^{\infty}\dfrac{q^{3j}}{(q;q)_{j}}\sum\limits_{n=0}^{\infty}\dfrac{(-1)^{n}\chi(n)q^{\binom{n+1}{2}+nj}}{(q;q)_{n}}.$
@@
$\varepsilon(q)$ $=(q;q)_{\infty}\sum\limits_{n=0}^{\infty}\dfrac{(-1)^{n}\chi(n)q^{\binom{n+1}{2}}}{(q;q)_{n}(q^{n+3};q)_{\infty}}$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}(1-q^{n+1})(1-q^{n+2})$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}(1-q^{n+1}(1+q)+q^{(n+1)+(n+2)})$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}-(1+q)\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+2}{2}}+\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+3}{2}}$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}+(1+q)\sum\limits_{n=1}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}}$
$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-2)q^{\binom{n+1}{2}}$
$=(2+q)+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}-2q(1+q)+(1+q)\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}}$
$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-2)q^{\binom{n+1}{2}}$
$=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\left(\chi(n)+\chi(n-1)+\chi(n-2)\right)q^{\binom{n+1}{2}}$
$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}+1}.$
Now, $\chi(n)+\chi(n-1)+\chi(n-2)=0$ for all $n$. Hence,
$\varepsilon(q)=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}+1}$
@@ 4. Concluding Remarks
It would be very interesting to provide a bijective proof of Theorem 1.6. The same procedure in the proof of Theorem 1.5 in Section 3 can be followed for $E_{m}(n)$ with $m\geq 4$; however, the outcome will be much messier.

### v2

# On Glaisher’s Partition Theorem

@@ Abstract.
Glaisher’s theorem states that the number of partitions of $n$ into parts which repeat at most m1m-1 times is equal to the number of partitions of $n$ into parts which are not divisible by $m$. The $m=2$ case is Euler’s famous partition theorem. Recently, Andrews, Kumar, and Yee gave two new partition functions $C(n)$ and $D(n)$ related to Euler’s theorem. Lin and Zhang extended their result to Glaisher’s theorem by generalizing $C(n)$. We generalize $D(n)$, prove an analogous partition identity for the $m=3$ case, and show that the general case is an example of an almost partition identity. We also provide a new series equal to Glaisher’s product both in the finite and infinite cases.
@@
Partitions, Partition Identities, Euler’s Theorem, Glaisher’s Theorem, $q$-series, Almost Partition Identities.
@@ 1. Introduction
([4, Theorem $1.1$]) For any positive integer $n$, we have
@@
Let $m\geq 2$ be an integer. Define $A_{m}(n)$ to be the number of partitions of n into parts repeating less than $m$ times and $B_{m}(n)$ to be the number of partitions of $n$ into parts not divisible by $m$. In 1883, Glaisher [5] obtained the following partition theorem which is a natural extension of Euler’s theorem (1.1).
@@
([6, Theorem $1.5$]) For any positive integer $n$, we have
@@
Next, for non-negative integers $a$, $b$, we define the $q$-binomial coefficient as
$\left[\begin{matrix}a+b\\ b\end{matrix}\right]_{q}:=\Bigg\{\begin{array}[]{lr}\dfrac{(q)_{a+b}}{(q)_{a}(q)_{b}}\quad\text{for }a,b\geq 0,\\ 0\qquad\qquad\text{otherwise}.\end{array}$
@@
(1.7) $\varepsilon_{m}(q):=\sum\limits_{n=0}^{\infty}E_{m}(n)q^{n}=\sum\limits_{n=0}^{\infty}q^{mn}(q^{n+1};q)_{\infty}\sum\limits_{j=1}^{m-1}(\zeta^{j}_{m}q^{n+1};q)_{\infty}.$
@@
(1.8) $\varepsilon_{3}(q):=\sum\limits_{n=0}^{\infty}E_{3}(n)q^{n}=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}+1},$
where $\chi_{3}(n)$ is defined as
$\chi_{3}(n):=\omega^{n}+\omega^{-n}=\Bigg\{\begin{array}[]{lr}2\qquad\,\,\,\text{if }3\mid n,\\ -1\quad\,\,\,\,\text{if }3\nmid n.\end{array}$
@@
Even more strikingly, we obtain another example of an almost partition identity (they are true for almost all $n$ but fail for an infinity of values; see [3]) for the general case which is as follows.
Theorem 1.7.
The identity
(1.10) $C_{m}(n)=\frac{1}{m}D_{m}(n)$
is true for almost all $n$. Moreover, if $N(x)$ is the number of times that (1.10) is true for $n<x$, then
$\lim\limits_{x\rightarrow\infty}\frac{N(x)}{x}=1.$
@@
In view of the generating function of $B_{2}(n)$ in [4, $(2.1)$], we also obtain the following finite $q$-series identity.
@@
The rest of the paper is organized as follows. In Sections 2, 3, and 4, we provide $q$-series (generating function) proofs of Theorem 1.4, 1.5, and 1.7 respectively.
@@
(3.1) $\varepsilon_{3}(q)=(q;q)_{\infty}\sum\limits_{j=0}^{\infty}\dfrac{q^{3j}}{(q;q)_{j}}\left((\omega q^{j+1};q)_{\infty}+(\omega^{-1}q^{j+1};q)_{\infty}\right).$
Now, define $\chi_{3}(n)$ as follows
$\chi_{3}(n):=\omega^{n}+\omega^{-n}=\Bigg\{\begin{array}[]{lr}2\qquad\,\,\,\text{if }3\mid n,\\ -1\quad\,\,\,\,\text{if }3\nmid n.\end{array}$
@@
(3.2) $\varepsilon_{3}(q)=(q;q)_{\infty}\sum\limits_{j=0}^{\infty}\dfrac{q^{3j}}{(q;q)_{j}}\sum\limits_{n=0}^{\infty}\dfrac{(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}+nj}}{(q;q)_{n}}.$
@@
$\varepsilon_{3}(q)$ $=(q;q)_{\infty}\sum\limits_{n=0}^{\infty}\dfrac{(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}}{(q;q)_{n}(q^{n+3};q)_{\infty}}$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}(1-q^{n+1})(1-q^{n+2})$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}(1-q^{n+1}(1+q)+q^{(n+1)+(n+2)})$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}-(1+q)\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+2}{2}}+\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+3}{2}}$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}+(1+q)\sum\limits_{n=1}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}}$
$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-2)q^{\binom{n+1}{2}}$
$=(2+q)+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}-2q(1+q)+(1+q)\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}}$
$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-2)q^{\binom{n+1}{2}}$
$=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\left(\chi_{3}(n)+\chi_{3}(n-1)+\chi_{3}(n-2)\right)q^{\binom{n+1}{2}}$
$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}+1}.$
Now, $\chi_{3}(n)+\chi_{3}(n-1)+\chi_{3}(n-2)=0$ for all $n$. Hence,
$\varepsilon_{3}(q)=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}+1}$
@@
4. Proof of Theorem 1.7
Define $\chi_{m}(n)$ as follows
$\chi_{m}(n):=\sum\limits_{j=1}^{m-1}\zeta^{jn}_{m}=\Bigg\{\begin{array}[]{lr}m-1\,\,\,\text{if }m\mid n,\\ -1\quad\,\,\,\,\,\text{if }m\nmid n.\end{array}$
Using [1, p. 1919, ($2.2.6$)] with $t=-\zeta^{j}_{m}q^{n}$ in (1.7), we get
(4.1) $\varepsilon_{m}(q)$ $=(q;q)_{\infty}\sum\limits_{n=0}^{\infty}\dfrac{q^{mn}}{(q)_{n}}\sum\limits_{k=0}^{\infty}\dfrac{(-1)^{k}\chi_{m}(k)q^{\binom{k+1}{2}+nk}}{(q;q)_{k}}.$
Using Euler’s identity [1, p. 1919, ($2.2.5$)] with $t=q^{k+m}$ after interchanging summations in (4.1), we then have
$\varepsilon_{m}(q)$ $=(q;q)_{\infty}\sum\limits_{k=0}^{\infty}\dfrac{(-1)^{k}\chi_{m}(k)q^{\binom{k+1}{2}}}{(q;q)_{k}(q^{k+m};q)_{\infty}}$
(4.2) $=\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k)q^{\binom{k+1}{2}}(q^{k+1};q)_{m-1}.$
Using [1, p. $36$, ($3.3.6$)] with the substitutions N=m1N=m-1 and $z=q^{k+1}$, we get
$\varepsilon_{m}(q)$ $=\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k)q^{\binom{k+1}{2}}\sum\limits_{j=0}^{m-1}(-1)^{j}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}q^{\binom{j+1}{2}+kj}$
$=\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k)\sum\limits_{j=0}^{m-1}(-1)^{j}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}q^{\binom{j+k+1}{2}}$
$=\sum\limits_{j=0}^{m-1}(-1)^{j}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k)q^{\binom{j+k+1}{2}}$
$=\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=j}^{\infty}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
$=-\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{j-1}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
$\qquad+\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
$=-\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{j-1}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
$\qquad+\sum\limits_{k=0}^{\infty}(-1)^{k}q^{\binom{k+1}{2}}\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}$
$=-\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{j-1}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
$\qquad+\sum\limits_{k=0}^{\infty}(-1)^{k}q^{\binom{k+1}{2}}\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)\left(\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}-1\right),$
where we have used the fact that $\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)=0$ in the second sum on the last line above. Hence, we have
(4.3) $\varepsilon_{m}(q)=P_{m}(q)+\sum\limits_{k=0}^{\infty}(-1)^{k}q^{\binom{k+1}{2}}\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)\left(\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}-1\right),$
where $P_{m}(q):=-\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{j-1}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$.
The polynomial
$\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)\left(\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}-1\right)$
in (4.3) above has at most $2^{m-1}-m$ terms. Now, there are at most $\sqrt{2x}+1$ triangular numbers less than $x$. Hence, the maximum possible number of non-zero coefficients in $\varepsilon_{m}(q)$ among the first $x$ terms is at most $(2^{m-1}-m)\cdot(\sqrt{2x}+1)$ leaving the number of coefficients that vanish as less than
$x-(2^{m-1}-m)\cdot(\sqrt{2x}+1)$
which is asymptotic to $x$ as $x\rightarrow\infty$. This shows that $E_{m}(n)$ is almost always zero which completes the proof of Theorem 1.7.
@@ 5. Concluding Remarks
It would be very interesting to provide a bijective proof of Theorem 1.6.